The current gain of a common emmiter amplifier is 69 . If the emmiter current is 7.0 mA , what will be the collecter current ?

Grade:

1 Answers

Khimraj
5 years ago
\beta = I_{c}/I_{b} = 69
 
\alpha = \beta /(1+\beta ) = 69/70 = I_{c}/I_{e}
Ic = (69/70)*Ie = 6.9 mA
Hope it clears.